OpenStudy (anonymous):

A cones height is 3 times the radius of the base. The area of the base is 49pi. What is the volume?

OpenStudy (goldphenoix):

To find the volume of a cone, you have to use the formula: \[\frac{ 1 }{ 3 } (Height \times Base)\] The height is 3 times the radius of the base. Or 3r. The area of the base is 49 pi. We want to find the radius so that we can find the height. Since the base of the cone is a circle, you have to do divide it by pi and by 2. So the radius would be 24.5. If the radius is 24.5, then the hieght is 73.5. It's time to plug it in: \[\frac{ 1 }{ 3 } (73.5 \times 49\Pi)\] 73.5 x 49 is equal to 3601.5. Next you divide it by 3 to get 1200.5. Times it by pi to get: 3771.48 unit^3.

OpenStudy (goldphenoix):

Hope this help! :)

OpenStudy (anonymous):

Thank you so much